One mole nitrogen oxide, NO, reacts with oxygen gas to produce a brown gas. How many moles NO must react with excess oxygen to release 120kJ of

heat?


1
Expert's answer
2022-06-07T01:45:48-0400

2NO + O2 = 2NO2 ∆H= -114 kJ


2NO + O2 + 114 kJ = 2NO2

2mol NO — 114 kJ

X — 120 kJ

X= 2 * 120 / 114 = 2,1 mol NO


Answer: 2,1 mol NO